OpenStudy (anonymous):

Find all the real solutions of x^3 - x^2 - 3x - 1 =0

OpenStudy (cwrw238):

f(-1) = -1^3 - (-1)^2 -3(-1) - 1 = = -1 - 1 + 3 - 1 = 0 so one solution is x = -1 and (x + 1) is a factor of the function

OpenStudy (cwrw238):

Dividing: x^2 - 2x - 1 ------------------ x+1) x^3 - x^2 - 3x - 1 x^3 + x^2 -------- -2x^2 - 3x -2x^2 - 2x ---------- -x - 1 -x - 1 ------ so f(x) = (x +1)(x^2 - 2x -1 )

OpenStudy (cwrw238):

to get the other roots solve (x^2 - 2x - 1) = 0

OpenStudy (anonymous):

x^3 - x^2 - 3x - 1 =0 f(x)=x^3 - x^2 - 3x - 1 now you have to take a value of x such that f(x)=0; if x=-1 then f(x)=0;lets see f(-1)=(-1)^3-(-1)^2-3(-1)-1 =-1-1+3-1=-3+3=0 as for x=-1 f(x) is 0 so (x-(-1))=(x+1) is one of the factor of the equation so now you can write x^3 - x^2 - 3x - 1 =0; =>x^2(x+1)-2x(x+1)-1(x+1)=0; [to be sure you can back calculate it] =>(x+1)(x^2-2x-1)=0 so x+1=0 or x=-1 and x^2-2x-1=0 or x=[-2pm sqrt{(-2)^2-4*1*(-1)}\2*1] [pm means plus minus] or x=1+sqrt (2 )or x= 1-sqrt (2) ans: x=-1; 1+sqrt (2 ); x= 1-sqrt (2)